OpenStudy (anonymous):

If the sequence is geometric, what is the common ratio? 500, 100, 20, 4,... 5 125 1/5 not geometric

OpenStudy (anonymous):

its 1/5

OpenStudy (anonymous):

1/5

OpenStudy (anonymous):

500/100=5, 100/20=5, 20/5=4

OpenStudy (anonymous):

1/5